OpenStudy (anonymous):

Integrate 2x+7/((x-3)(x-4))

OpenStudy (phi):

is the problem \[\int\limits_{}^{}\frac{2x+7}{(x-3)(x-4)}dx\] ? If so, do you know how to do partial fractions?

OpenStudy (anonymous):

Yes ands no

OpenStudy (phi):

Try to do the partial fraction expansion. \[\frac{2x+7}{(x-3)(x-4)}=\frac{A}{(x-3)}+\frac{B}{(x-4)}\]

OpenStudy (anonymous):

I need help

OpenStudy (phi):

for the partial fraction expansion, add the 2 fractions. for the first fraction with A on top, you multiply top and bottom by (x-4) (to get the common denominator).

OpenStudy (phi):

for the 2nd fraction, multiply top and bottom by (x-3) now add.